# Comparing Task Graph Scheduling Algorithms: An Adversarial Approach

### v2

@@
Abstract.
Scheduling a task graph representing an application over a heterogeneous network of computers is a fundamental problem in distributed computing. It is known to be not only NP-hard but also not polynomial-time approximable within a constant factor. As a result, many heuristic algorithms have been proposed over the past few decades. Yet it remains largely unclear how these algorithms compare to each other in terms of the quality of schedules they produce. We identify gaps in the traditional benchmarking approach to comparing task scheduling algorithms and propose a simulated annealing-based adversarial analysis approach called PISA to help address them. We also introduce SAGA, a new open-source library for comparing task scheduling algorithms. We use SAGA to benchmark 15 algorithms on 16 datasets and PISA to compare the algorithms in a pairwise manner. Algorithms that appear to perform similarly on benchmarking datasets are shown to perform very differently on adversarially chosen problem instances. Interestingly, the results indicate that this is true even when the adversarial search is constrained to selecting among well-structured, application-specific problem instances. This work represents an important step towards a more general understanding of the performance boundaries between task scheduling algorithms on different families of problem instances.
@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM Symposium on Principles of Distributed Computing; June 17–21, 2024; Nantes, Franceprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Task graph scheduling is a fundamental problem in distributed computing. Essentially, the goal is to assign computational tasks to different compute nodes in such a way that minimizes/maximizes some performance metric (e.g., total execution time, energy consumption, throughput, etc.). In this paper, we will focus on the task scheduling problem concerning heterogeneous task graphs and compute networks with the objective of minimizing makespan (total execution time) under the related machines model [^1] In the related machines model, if the same task executes faster one some compute node $n_{1}$ than on node $n_{2}$, then $n_{1}$ must execute all tasks faster than $n_{2}$ ($n_{1}$ is strictly faster than $n_{2}$). Observe this model cannot describe multi-modal distributed systems, where certain classes of tasks (e.g., GPU-heavy tasks) might run better/worse on different types of machines (e.g., those with or without GPUs). The related machines model as it pertains to the task scheduling problem we study in this paper is described further in Section 1.1.. As this problem is NP-hard (Garey and Johnson, 1979) and also not polynomial-time approximable within a constant factor (Bazzi and Norouzi-Fard, 2015), many heuristic algorithms have been proposed. Despite the overwhelming number of algorithms now available in the literature, open-source implementations of them are scarce. Even worse, the datasets on which they were evaluated on on are often not available and the implementations that do exist are not compatible with each other (different programming languages, frameworks, problem instance formats, etc.). In order to address this technological shortcoming and enable this study, we built SAGA, a Python framework for running, evaluating, and comparing task scheduling algorithms.
@@ 1. Introduction
  1. (1)
    Introduces SAGA (Authors, 2023b), an open-source, modular, and extensible Python package with implementations of many well-known task scheduling algorithms and tools for benchmarking and adversarial analysis.
  2. (2)
@@
  3. (3)
    Proposes PISA, a novel simulated annealing-based adversarial analysis method for comparing algorithms that identifies problem instances for which a given algorithm maximally underperforms another.
  4. (4)
@@
The rest of the paper is organized as follows. We formally define the problem in Section 1.1. In Section 2, we provide a brief history of the task scheduling problem and related work. In Section 3, we introduce SAGA, a Python library for running, evaluating, and comparing task scheduling algorithms. Then, in Section 4, we present the results of benchmarking 15 algorithms on 16 datasets. In Section 5, we present the main contribution of this paper: PISA, a simulated annealing-based adversarial analysis method for comparing task scheduling algorithms. In Section 5.1 we present the results of comparing the 15 algorithms in a pairwise manner using this method. In Section 6, we demonstrate how our proposed technique can be tailored to application-specific scenarios where certain properties of the task graph and/or network are known ahead of time. We conclude the paper in Section 7 with a short discussion on the implications of this work and directions for future research.
1.1. Problem Definition
@@
Let us denote the task graph as $G=(T,D)$, where $T$ is the set of tasks and $D$ contains the directed edges or dependencies between these tasks. An edge $(t,t^{\prime})\in D$ implies that the output from task $t$ is required input for task $t^{\prime}$. Thus, task $t^{\prime}$ cannot start executing until it has received the output of task $t$. This is often referred to as a precedence constraint. For a given task $t\in T$, its compute cost is represented by $c(t)\in\mathbb{R}^{+}$ and the size of the data exchanged between two dependent tasks, $(t,t^{\prime})\in D$, is $c(t,t^{\prime})\in\mathbb{R}^{+}$. Let $N=(V,E)$ denote the compute node network, where $N$ is a complete undirected graph. $V$ is the set of nodes and $E$ is the set of edges. The compute speed of a node $v\in V$ is $s(v)\in\mathbb{R}^{+}$ and the communication strength between nodes $(v,v^{\prime})\in E$ is $s(v,v^{\prime})\in\mathbb{R}^{+}$. Under the related machines model (Graham, 1969), the execution time of a task $t\in T$ on a node $v\in V$ is $\frac{c(t)}{s(v)}$, and the data communication time between tasks $(t,t^{\prime})\in D$ from node $v$ to node $v^{\prime}$ (i.e., $t$ executes on $v$ and $t^{\prime}$ executes on $v^{\prime}$) is $\frac{c(t,t^{\prime})}{s(v,v^{\prime})}$.
@@
The goal is to schedule the tasks on different compute nodes in such a way that minimizes the makespan (total execution time) of the task graph. Let $\mathcal{A}$ denote a task scheduling algorithm. Given a problem instance $(N,G)$ which represents a network/task graph pair, let $S_{\mathcal{A},N,G}$ denote the schedule produced by $\mathcal{A}$ for $(N,G)$. A schedule is a mapping from each task to a triple $(v,r,e)$ where $v$ is the node on which the task is scheduled, $r$ is the start time, and $e$ is the end time. A valid schedule must satisfy the following properties
  • All tasks must be scheduled: for all $t\in T$, $S_{\mathcal{A},N,G}(t)=(v,r,e)$ must exist such that $v\in V$ and $0\leq r\leq e$.
  • All tasks must have valid start and end times:
    $\forall t\in T,S_{\mathcal{A},N,G}(t)=(v,r,e)\implies e-r=\frac{c(t)}{s(v)}$
  • Only one task can be scheduled on a node at a time (i.e., their start/end times cannot overlap):
    $\forall t,t^{\prime}\in T,t\neq t^{\prime},S_{\mathcal{A},N,G}(t)=(v,r,e)\land S_{\mathcal{A},N,G}(t^{\prime})=(v,r^{\prime},e^{\prime})\implies e\leq r^{\prime}\lor e^{\prime}\leq r$
@@
    $\forall(t,t^{\prime})\in D,S_{\mathcal{A},N,G}(t)=(v,r,e)\land S_{\mathcal{A},N,G}(t^{\prime})=(v^{\prime},r^{\prime},e^{\prime})\implies e+\frac{c(t,t^{\prime})}{s(v,v^{\prime})}\leq r^{\prime}$
@@
Figure 1. Example problem instance and schedule.
@@
Figure 1 depicts an example problem instance (task graph and network) and solution (schedule). We define the makespan of the schedule $S_{\mathcal{A},N,G}$ as the time at which the last task finishes executing:
$M_{\mathcal{A}(N,G)}=\max_{t\in T\mid S_{\mathcal{A},N,G}(t)=(v,r,e)}e$
@@
Because the problem of minimizing makespan is NP-hard for this model (Garey and Johnson, 1979), many heuristic algorithms have been proposed. Traditionally, these heuristic algorithms are evaluated on a set of problem instances and compared against other algorithms based on their makespan ratio, which for a given problem instance $(N,G)$ is the makespan of the schedule produced by the algorithm divided by the minimum makespan of the schedules produced by the baseline algorithms. The makespan ratio of an algorithm $\mathcal{A}$ against a set of baseline algorithms $\mathcal{A}_{1},\mathcal{A}_{2},\ldots$ for a problem instance $(N,G)$ can be written
$\frac{M_{\mathcal{A}(N,G)}}{\min\left(M_{\mathcal{A}_{1}(N,G)},M_{\mathcal{A}_{2}(N,G)},M_{\mathcal{A}_{3}(N,G)},\ldots\right)}.$
@@
Makespan ratio is a commonly used concept in the task scheduling literature (Topcuoglu et al., 1999; Wang and Sinnen, 2018). It is common to collect the makespan ratios of an algorithm (against a set of baseline algorithms) for a dataset of problem instances. System designers use this technique, called benchmarking, to decide which algorithm(s) best support(s) their application. In this paper, we highlight the shortcomings of this approach by discussing examples of (and providing methods for the automatic discovery of) in-family problem instances on which algorithms that appear to perform well on the original dataset perform very poorly.
2. Related Work
@@
A comprehensive survey by Ronald Graham in 1979 classified multiple variants of the task scheduling problem and proposed a structured approach to analyze their complexity classes (Graham et al., 1979). Though the variant of task scheduling discussed in this paper wasn’t directly addressed in their study, the established reduction framework paved the way for its NP-completeness determination. The survey addressed many facets of task graph scheduling, such as processor heterogeneity, precedence constraints, and makespan minimization, but did not consider inter-processor communication. This gap was addressed a decade later, in 1989, by Hwang et al., who introduced the concept of task graph scheduling with inter-processor communication (albeit only considering homogeneous processors). They proposed the ETF (Earliest Task First) algorithm and proved that it produces schedules with makespan of at most $(2-1/n)\omega_{\text{opt}}^{(i)}+C$ where $\omega_{\text{opt}}^{(i)}$ is the optimal makespan without considering inter-processor communication delays and $C$ is the worst-case communication requirements over a terminal chain of tasks (one which determines the makespan).
@@
In other words, list scheduling algorithms first decide on a valid topological sort of the task graph, and then use it to schedule tasks greedily according to some objective function (ETF, mentioned above, is a list scheduling algorithm). Two of the most well-known heuristic algorithms for task graph scheduling, HEFT (Heterogeneous Earliest Finish Time) and CPoP (Critical Path on Processor) (Topcuoglu et al., 1999), are list scheduling algorithms. Many other heuristic algorithms have been proposed and experimentally evaluated on different datasets of task graphs and networks (see benchmarking/comparison/survey papers (Braun et al., 1999a, 2001; Canon et al., 2008; Kwok and Ahmad, 1998; Wang and Sinnen, 2018)). Other paradigms studied in the literature are cluster-scheduling (Wang and Sinnen, 2018) and meta-heuristic (Houssein et al., 2021) algorithms. Cluster scheduling involves dividing the task graph into groups of tasks to execute on a single node. Meta-heuristic approaches (e.g., simulated annealing or genetic algorithms) have been shown to work well in some situations, but generally take longer to run and can be difficult to tune (Braun et al., 1999b).
In this paper, we propose a simulated annealing-based (Kirkpatrick et al., 1983) approach to finding problem instances on which an algorithm performs maximally poorly compared to a given baseline algorithm. This approach is inspired by the recent work of Namyar et al. who use simulated annealing and other techniques to find adversarial instances for heuristics that solve convex optimization problems (Namyar et al., 2022). To the best of our knowledge, this is the first work to propose an automated method for adversarial analysis of task scheduling algorithms.
3. The SAGA Framework
We built SAGA to overcome the notable lack of publicly available datasets and task scheduling algorithm implementations. SAGA is a Python library for running, evaluating, and comparing task scheduling algorithms. It currently contains implementations of 17 algorithms using a common interface. It also includes interfaces for generating, saving, and loading datasets for benchmarking. Finally, it includes an implementation of the main contribution of this paper: PISA, a simulated annealing-based adversarial analysis met hod for comparing algorithms. For more information on SAGA, we refer the reader to the technical report (Authors, 2023b).
Table 1 lists the 17 algorithms currently implemented in SAGA.
Table 1. Schedulers implemented in SAGA
@@ 3. The SAGA Framework
BIL Best Imaginary Level (Oh and Ha, 1996)
@@
CPoP Critical Path on Processor (Topcuoglu et al., 1999)
Duplex Duplex (Braun et al., 2001)
ETF Earliest Task First (Hwang et al., 1989)
@@
FCP Fast Critical Path (Radulescu and van Gemund, 2000)
FLB Fast Load Balancing (Radulescu and van Gemund, 2000)
GDL Generalized Dynamic Level (Sih and Lee, 1993b)
HEFT Heterogeneous Earliest Finish Time (Topcuoglu et al., 1999)
MaxMin MaxMin (Braun et al., 2001)
MCT Minimum Completion Time (Armstrong et al., 1998)
MET Minimum Execution Time (Armstrong et al., 1998)
MinMin MinMin (Braun et al., 2001)
OLB Opportunistic Load Balancing (Armstrong et al., 1998)
@@
WBA Workflow-Based application (Blythe et al., 2005)
To orient the reader, we provide a brief description of each of these algorithms along with their scheduling complexity, the model they were designed for, performance guarantees (if any), datasets they have been evaluated on, and other algorithms they have been evaluated against in Appendix A.1. Because the BruteForce and SMT scheduling algorithms take much longer to run (exponential time) than the other algorithms, they are not included in the benchmarking or adversarial analysis results reported in this paper.
SAGA also includes a set of tools for generating, saving, and loading datasets. Table 2 lists the 16 dataset generators currently included in SAGA.
Table 2. Datasets available in SAGA

### v3

# Los Angeles, USA

@@
Abstract
Scheduling a task graph representing an application over a heterogeneous network of computers is a fundamental problem in distributed computing. It is known to be not only NP-hard but also not polynomial-time approximable within a constant factor. As a result, many heuristic algorithms have been proposed over the past few decades. Yet it remains largely unclear how these algorithms compare to each other in terms of the quality of schedules they produce. We identify gaps in the traditional benchmarking approach to comparing task scheduling algorithms and propose a simulated annealing-based adversarial analysis approach called PISA to address them. We also introduce SAGA, a new open-source library for comparing task scheduling algorithms. We use SAGA to benchmark 15 algorithms on 16 datasets and PISA to compare the algorithms in a pairwise manner. Algorithms that appear to perform similarly on benchmarking datasets are shown to perform very differently on adversarially chosen problem instances. Interestingly, the results indicate that this is true even when the adversarial search is constrained to selecting among well-structured, application-specific problem instances. We present the first known lower bounds on the performance of many of the algorithms considered in this paper compared to other popular scheduling algorithms. This work represents an important step towards a more general understanding of the performance boundaries between task scheduling algorithms on different families of problem instances.
Index Terms:
@@ Abstract
I Introduction
Task graph scheduling is a fundamental problem in distributed computing. It plays a crucial role in optimizing performance across various domains, including IoT systems, scientific workflows, and load balancing in distributed networks. Essentially, the goal is to assign computational tasks to different compute nodes in such a way that minimizes/maximizes some performance metric (e.g., total execution time, energy consumption, throughput, etc.). In this paper, we will focus on the offline (also called static or compile-time) task scheduling problem concerning heterogeneous task graphs and compute networks with the objective of minimizing makespan (total execution time) under the related machines model [^1] In the related machines model, if the same task executes faster on some compute node $n_{1}$ than on node $n_{2}$, then $n_{1}$ must execute all tasks faster than $n_{2}$ ($n_{1}$ is strictly faster than $n_{2}$). Observe this model cannot describe multi-modal distributed systems, where certain classes of tasks (e.g., GPU-heavy tasks) might run better/worse on different types of machines (e.g., those with or without GPUs). The related machines model as it pertains to the task scheduling problem we study in this paper is described further in Section II.. As this problem is NP-hard [1] and also not polynomial-time approximable within a constant factor [2], many heuristic algorithms have been proposed. Despite the overwhelming number of algorithms now available in the literature, open-source implementations of them are scarce. Even worse, the datasets on which they were evaluated are often not available and the implementations that do exist are not compatible with each other (different programming languages, frameworks, problem instance formats, etc.). In order to address this technological shortcoming and enable this study, we built SAGA [3], a Python framework for running, evaluating, and comparing task scheduling algorithms.
@@ I Introduction
  1. 1.
    Introduces SAGA, an open-source, modular, and extensible Python package with implementations of many well-known task scheduling algorithms and tools for benchmarking and adversarial analysis.
  2. 2.
@@
  3. 3.
    Proposes PISA, a novel simulated annealing-based adversarial analysis method for comparing algorithms that identifies problem instances for which a given algorithm maximally under-performs another.
  4. 4.
@@
The rest of the paper is organized as follows. We formally define the problem in Section II. In Section III, we provide a brief history of the task scheduling problem and related work. In Section IV, we introduce SAGA, a Python library for running, evaluating, and comparing task scheduling algorithms. Then, in Section V, we present the results of benchmarking 15 algorithms on 16 datasets. In Section VI, we present the main contribution of this paper: PISA, a simulated annealing-based adversarial analysis method for comparing task scheduling algorithms. In Section VI-A we present the results of comparing the 15 algorithms in a pairwise manner using this method. We take a closer look at one example comparing HEFT and CPoP (two well-known task scheduling algorithms [4]) in Section VI-B, demonstrating how this framework might be used to discover the reasons why an algorithm performs well on some problem instances and poorly on others. In Section VII, we demonstrate how our proposed technique can be tailored to application-specific scenarios where certain properties of the task graph and/or network are known ahead of time. In Section VII, we show that this approach is not limited to finding only pathological (perhaps unrealistic) problem instances, but can also be used to find adversarial instances that are well-structured and application-specific. We conclude the paper in Section VIII with a short discussion on the implications of this work and directions for future research.
II Problem Definition
@@
Let us denote the task graph as $G=(T,D)$, where $T$ is the set of tasks and $D$ contains the directed edges or dependencies between them. An edge $(t,t^{\prime})\in D$ implies that the output from task $t$ is required input for task $t^{\prime}$. Thus, task $t^{\prime}$ cannot start executing until it has received the output of task $t$. This is often referred to as a precedence constraint. The compute cost of given task $t\in T$ is represented by $c(t)\in\mathbb{R}^{+}$ and the size of the data exchanged between two inter-dependent tasks, $(t,t^{\prime})\in D$, is $c(t,t^{\prime})\in\mathbb{R}^{+}$. Let $N=(V,E)$, a complete undirected graph, denote the compute node network. $V$ is the set of nodes and $E$ is the set of edges. The compute speed of a node $v\in V$ is $s(v)\in\mathbb{R}^{+}$ and the communication strength between nodes $(v,v^{\prime})\in E$ is $s(v,v^{\prime})\in\mathbb{R}^{+}$. Under the related machines model [5], the execution time of a task $t\in T$ on a node $v\in V$ is $\frac{c(t)}{s(v)}$, and the data communication time between tasks $(t,t^{\prime})\in D$ from node $v$ to node $v^{\prime}$ (i.e., $t$ executes on $v$ and $t^{\prime}$ executes on $v^{\prime}$) is $\frac{c(t,t^{\prime})}{s(v,v^{\prime})}$.
@@
The goal is to schedule the tasks on different compute nodes in such a way that minimizes the makespan (total execution time) of the task graph. For a given problem instance $(N,G)$ which represents a network/task graph pair, a schedule is a set of tuples of the form $(t,v,r)$ where $t\in T$ is the scheduled task, $v\in V$ is the node on which task $t$ is scheduled to run, and $r\in\mathbb{R}^{+}$ is the start time. A valid schedule $S$ must satisfy the following properties
  • All tasks must be scheduled exactly once:
    $\forall t\in T,\exists(t,v,r)\in S$
    $\forall(t,v,r),(t^{\prime},v^{\prime},r^{\prime})\in S,t=t^{\prime}\Rightarrow(v,r)=(v^{\prime},r^{\prime})$
@@
    $\forall(t,v,r),(t^{\prime},v^{\prime},r^{\prime})$ $\in S,(t,t^{\prime})\in D$
    $\Rightarrow r+\frac{c(t)}{s(v)}+\frac{c(t,t^{\prime})}{s(v,v^{\prime})}\leq r^{\prime}$
@@
Figure 1: Example problem instance and schedule.
@@
Figure 1 depicts an example problem instance (task graph and network) and solution (schedule) as a Gantt chart. We define the makespan of a schedule $S$ as the time at which the last task finishes executing:
$m(S)=\max_{(t,v,r)\in S}r+\frac{c(t)}{s(v)}$
@@
Because the problem of minimizing makespan is NP-hard for this model [1], many heuristic algorithms have been proposed. Traditionally, these heuristic algorithms are evaluated on a set of problem instances and compared against other algorithms based on their makespan ratio, which for a given problem instance $(N,G)$ is the makespan of the schedule produced by the algorithm divided by the minimum makespan of the schedules produced by the baseline algorithms. Let $S_{\mathcal{A},N,G}$ denote the schedule produced by algorithm $\mathcal{A}$ on problem instance $(N,G)$. Then the makespan ratio of an algorithm $\mathcal{A}$ against a set of baseline algorithms $\mathcal{A}_{1},\mathcal{A}_{2},\ldots$ for a problem instance $(N,G)$ can be written
$\frac{m\left(S_{\mathcal{A},N,G}\right)}{\min\left\{m\left(S_{\mathcal{A}_{1},N,G}\right),m\left(S_{\mathcal{A}_{2},N,G}\right),m\left(S_{\mathcal{A}_{3},N,G}\right),\ldots\right\}}.$
@@
Makespan ratio is a well-known concept in the task scheduling literature (sometimes referred to by other names such as “schedule length ratio”) [4, 6]. It is common to collect the makespan ratios of an algorithm against a set of baseline algorithms for a dataset of problem instances. System designers use this technique, called benchmarking, to decide which algorithm(s) best support(s) their application. We highlight the shortcomings of this approach by presenting examples (and providing methods for the automatic discovery) of in-family problem instances on which algorithms that appear to perform well on the original dataset perform very poorly.
III Related Work
@@
A comprehensive survey by Ronald Graham in 1979 classified multiple variants of the task scheduling problem and proposed a structured approach to analyze their complexity classes [7]. Though the variant of task scheduling discussed in this paper wasn’t directly addressed in their study, the established reduction framework paved the way for its NP-completeness determination. The survey addressed many facets of task graph scheduling, such as processor heterogeneity, precedence constraints, and makespan minimization, but did not consider inter-processor communication. This gap was addressed a decade later, in 1989, by Hwang et al., who introduced the concept of task graph scheduling with inter-processor communication (albeit only considering homogeneous processors) [8]. They proposed the ETF (Earliest Task First) algorithm and proved that it produces schedules with makespan of at most $(2-1/n)\omega_{\text{opt}}^{(i)}+C$ where $\omega_{\text{opt}}^{(i)}$ is the optimal makespan without considering inter-processor communication delays and $C$ is the worst-case communication requirement over a terminal chain of tasks (one which determines the makespan).
@@
In other words, list scheduling algorithms first decide on a valid topological sort of the task graph, and then use it to schedule tasks greedily according to some objective function. ETF, mentioned above, is a list scheduling algorithm. Two of the most well-known heuristic algorithms for task graph scheduling, HEFT (Heterogeneous Earliest Finish Time) and CPoP (Critical Path on Processor) [4], are also list scheduling algorithms. Many other heuristic algorithms have been proposed and experimentally evaluated on different datasets of task graphs and networks (see benchmarking/comparison/survey papers [9, 10, 11, 12, 6]). Other paradigms studied in the literature are cluster-scheduling [6] and meta-heuristic [13] algorithms. Cluster scheduling involves dividing the task graph into groups of tasks to execute on a single node. Meta-heuristic approaches (e.g., simulated annealing or genetic algorithms) have been shown to work well in some situations, but generally take longer to run and can be difficult to tune [14].
In this paper, we propose a simulated annealing-based [15] approach to find problem instances on which an algorithm performs maximally poorly compared to a given baseline algorithm. This approach is inspired by the recent work of Namyar et al. who use simulated annealing and other techniques to find adversarial instances for heuristics that solve convex optimization problems [16]. To the best of our knowledge, this is the first work to propose an automated method for adversarial analysis of task scheduling algorithms.
IV The SAGA Framework
We built SAGA to overcome the notable lack of publicly available datasets and task scheduling algorithm implementations. SAGA is a Python library for running, evaluating, and comparing task scheduling algorithms. It currently contains implementations of 17 algorithms using a common interface. It also includes interfaces for generating, saving, and loading datasets for benchmarking. Finally, it includes an implementation of the main contribution of this paper: PISA, a simulated annealing-based adversarial analysis method for comparing algorithms. For more information on SAGA, we refer the reader to the open-source implementation [3]. Table I lists the 17 algorithms currently implemented in SAGA.
TABLE I: Schedulers implemented in SAGA
@@ IV The SAGA Framework
BIL Best Imaginary Level [17]
@@
CPoP Critical Path on Processor [4]
Duplex Duplex [10]
ETF Earliest Task First [8]
@@
FCP Fast Critical Path [18]
FLB Fast Load Balancing [18]
GDL Generalized Dynamic Level [19]
HEFT Heterogeneous Earliest Finish Time [4]
MaxMin MaxMin [10]
MCT Minimum Completion Time [20]
MET Minimum Execution Time [20]
MinMin MinMin [10]
OLB Opportunistic Load Balancing [20]
@@
WBA Workflow-Based application [21]
To orient the reader, we provide a brief description of each of these algorithms along with their scheduling complexity, the model they were designed for, performance guarantees (if any), datasets they have been evaluated on, and other algorithms they have been evaluated against.
IV-A Scheduling Algorithm Descriptions
To orient the reader, we provide a brief description of each of the algorithms implemented in SAGA.
CPoP (Critical Path on Processor) was proposed in the same paper as HEFT (Heterogeneous Earliest Finish Time). Both are list scheduling algorithms with scheduling complexity $O\left(|T|^{2}|V|\right)$. No formal bounds for HEFT and CPoP are known. HEFT works by first prioritizing tasks based on their upward rank, which is essentially the length of the longest chain of tasks (considering average execution and communication times). Then, it greedily schedules tasks in this order to the node that minimizes the task completion time given previously scheduled tasks. CPoP is similar but uses a slightly different priority metric. The biggest difference between the two is that CPoP always schedules critical path tasks (those on the longest chain in the task graph) to execute on the fastest node. Both algorithms were evaluated on random graphs (the process for generating graphs is described in the paper) and on real task graphs for Gaussian Elimination and FFT applications. They were compared against different scheduling algorithms: Mapping Heuristic (similar to HEFT without insertion) [22], Dynamic-Level Scheduling [23], and Levelized Min Time [^2] We could not find the original paper that proposes this algorithm.. Schedule Length Ratios (makespan scaled by the sum of minimum computation costs of all tasks), speedup, and schedule generation times are reported.
The MinMin, MaxMin, and Duplex list scheduling algorithms may have been proposed many times independently, but exceptionally clear definitions can be found in a paper that compares them to OLB, MET, MCT, and a few meta-heuristic algorithms (e.g., genetic algorithms and simulated annealing) [10]. MinMin schedules tasks by iteratively selecting the task with the smallest minimum completion time (given previously scheduled tasks) and assigning it to the corresponding node. MaxMin, on the other hand, schedules tasks by iteratively selecting the task with the largest minimum completion time and assigning it to the the corresponding node. Duplex simply runs both MinMin and MaxMin and returns the schedule with the smallest makespan. In the paper, the authors evaluate these algorithms on independent non-communicating tasks with uniformly random costs and heterogeneous compute nodes with uniformly random speeds. MinMin (and therefore also Duplex) is shown to generate schedules with low makespan compared to the other algorithms while relatively high makespans for MaxMin are reported.
FCP (Fast Critical Path) and FLB (Fast Load Balancing) both have a runtime of $O(|T|\log\left(|V|)+|D|\right)$ and were designed for heterogeneous task graphs and heterogeneous node speeds but homogeneous communication strengths. The algorithms were evaluated on three types of task graphs with different structures based on real applications (LU decomposition, Laplace equation solver, and a stencil algorithm) with varied CCR and uniformly random task costs. Both algorithms were shown to perform well compared to HEFT and ERT [24] despite their lower schedule generation times.
GDL (Generalized Dynamic Level), also called DLS (Dynamic Level Scheduling), is a variation on list scheduling where task priorities are updated each time a task is scheduled. Due to this added computation in each iteration, the complexity of DLS is $O(|V|^{3}|T|)$ (a factor $|V|$ greater than HEFT and CPoP). GDL was originally designed for the very general unrelated machines model and was shown to outperform HDLFET [19] on randomly generated problem instances (though the method used for generating random task graphs is not well-described) and on four real digital signal processing applications (two sound synthesis algorithms, a telephone channel simulator, and a quadrature amplitude modulation transmitter).
OLB (Opportunistic Load Balancing) has a runtime of just $O(|T|)$ and was designed for independent tasks under the unrelated machines model. Probably useful only as a baseline for understanding the performance of other algorithms, OLB schedules tasks in arbitrary order on the earliest available compute node. Its performance has been shown to be significantly worse than MET, MCT, and LBA [20].
Because the BruteForce and SMT scheduling algorithms take much longer to run (exponential time) than the other algorithms, they are not included in the benchmarking or adversarial analysis results reported in this paper.
IV-B Datasets
SAGA also includes a set of tools for generating, saving, and loading datasets. Table II lists the 16 dataset generators currently included in SAGA.
TABLE II: Datasets available in SAGA